1. The motion of a particle is defined by the relation
1
Expert's answer
2020-09-01T10:58:51-0400

The motion of a particle is defined by the relation "x=6t^{4}-2t^{3}-12t^{2}+3t+3", where x and t expressed in meters and seconds. Determine the time, the position and the velocity when a=0

"v=\frac{dx}{dt}(6t^{4}-2t^{3}-12t^{2}+3t+3)"

"=(4\times6)t^{3}-(3\times2)t^{2}-(12\times2)t+3"

"=24t^{3}-6t^{2}-24t+3 (1)"

"a=\frac{d}{dt}(v)=(24\times3)t^{2}-(2\times6)t-24"

"=72t^{2}-12t-24"

so "0=72t^{2}-12t-24"

"6t^{2}-t-2=0"

"t_{1}=-\frac{1}{2}, t_{2}=\frac{2}{3}"

taking the positive time as correct we substitute to (1)

so "v=24(\frac{2}{3})^{3}-6(\frac{2}{3})^{2}-24(\frac{2}{3})+3"

"=24(0.2962)-6(0.4444)-16+3"

"=7.1088-2.664-16+3=-8.5552 m\/sec"

Also "x-(6\times0.1975)-(2\times0.2962)-(12\times0.4444)+2+3"

"=1.185-0.5924-5.3328+2+3"

"=0.2598 m"